OpenStudy (anonymous):

please help me solve 3xsquared+8x+2=0 using the quadratic formula, I got almost to the end but the last of it is giving me some issue

OpenStudy (jlastino):

What seems to be the problem in your solution?

OpenStudy (anonymous):

(3x-1)^2 = 1/4 then (3x-1) = Sqrt(1/4) = 1/2 or -1/2 so 3x-1 = 1/2 or -1/2. Add one to both sides and get: 3x = 3/2 or 1/2 Divide by three on both sides to get: x = 1/2 or 1/6 Hope this helps.

OpenStudy (anonymous):

well thats a hint to solve

OpenStudy (anonymous):

I got to 

OpenStudy (anonymous):

but i have to use the quadratic formula

OpenStudy (jlastino):

You need help simplifying sqrt40? How do you represent 40 in terms of prime numbers?

OpenStudy (anonymous):